OpenStudy (anonymous):

Find the slope of the line joining (-3,2) and (3,-8)

OpenStudy (turingtest):

sorry, typo... The slope of the line between two points\[ (x_1,y_1)\]and\[(x_2,y_2) \]is given by \[m={y_2-y_1\over x_2-x_1}\]

OpenStudy (anonymous):

would it be -8-2 / 3-3 = -10/-6

OpenStudy (turingtest):

almost, but be careful with the minus sign:\[m={y_2-y_1\over x_2-x_1}={-8-2\over3-(-3)}={-10\over6}={-5\over3}\]

OpenStudy (anonymous):

thanks but how did we get -5/3

OpenStudy (anonymous):

divide by 2?

OpenStudy (turingtest):

yes, that is how the fraction simplifies

OpenStudy (anonymous):

ohh ok thanks so much

OpenStudy (turingtest):

like 2/4=1/2 for instance